OpenStudy (j.f):

Medal A floor that measures 10 feet by 20 feet is to be tiled with square tiles that measure 36 sq. in. each. How many tiles will be needed?

OpenStudy (linn99123):

isnt the fomula like length times width or something ? XD

OpenStudy (linn99123):

J.F here http://prntscr.com/6irb5l

OpenStudy (linn99123):

calculators on this website http://www.besttile.com/TileCalculator.aspx?name=Resource-Center

OpenStudy (j.f):

sorry I wasn't paying attention:/

OpenStudy (j.f):

I know how to get the total square footage of the whole area which is 200. So to convert to square inches I think it would be (200in^2)(12in/ft)^2= (200/144)=25/18= 1 7/18 sq. inch correct?

OpenStudy (j.f):

If so it would be linear Inches/piece length= quantity. 25/18/36= 25/648

OpenStudy (j.f):

nvm. I think it would be 25/18x36= 50 tiles